OpenStudy (anonymous):

The larger of two numbers is 3 times the smaller number. Four times the smaller number, decreased by 2 equals 6 more than the larger number. Find both numbers.

OpenStudy (anonymous):

so call the large number a, the small number b

OpenStudy (anonymous):

a is 3 times b so 3a = b

OpenStudy (anonymous):

4 times the smaller number (4b) decreased by 2 (4b - 2) equals 6 more than the larger number (6a) so 4b - 2 = 6a

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

I'm not quite sure how you would solve it from there. Do you get both the variables on one side?

OpenStudy (anonymous):

wel the first equation (3a) = b. so you know what b equals. 3a. so the second equation 4b - 2 = 6a, well you can just replace b with 3a, because b = 3a. so that would give you 4(3a) - 2 = 6a

OpenStudy (anonymous):

now you need to solve that to get the value of a

OpenStudy (anonymous):

So a would equal .3 and b would equal 1?

OpenStudy (anonymous):

yeah i get a as 1/3 too. and yeah then b would be 1